OpenStudy (anonymous):

Multiply. (x + 8)(x – 6)

OpenStudy (anonymous):

ANSWER CHOICES x2 + 2x – 48, x2 – 48, x2 – 2x – 48, x2 – 48x + 2,

OpenStudy (anonymous):

i dont know what that is

OpenStudy (anonymous):

Okay..

OpenStudy (anonymous):

Solve this: \[x(x-6) + 8(x-6)\] Distribute x and 8 inside the brackets..

OpenStudy (anonymous):

2x-6x+ 8x- 48 ?

OpenStudy (anonymous):

x multiplied by x will give you \(x^2\) and not 2x..

OpenStudy (anonymous):

rest is fine..

OpenStudy (anonymous):

And tell me what is: -6x + 8x = ???

OpenStudy (anonymous):

oh okay so x^2-6x+8x-48

OpenStudy (anonymous):

umm, 2x^2?

OpenStudy (anonymous):

you have to add them See the difference: 8x - 6x = 2x (in case of addition and subtraction) 8x*6x = 48\(x^2\) ( in case of multiplication) So now tell me: -6x + 8x = ??

OpenStudy (anonymous):

i still get 2 idont know

OpenStudy (anonymous):

-6x + 8x = 2x not 2 \(x^2\)

OpenStudy (anonymous):

What what is the final answer you got ??

OpenStudy (anonymous):

x^2-2x-48?